OpenStudy (anonymous):

Solve 3(4x − 2x) = 8

OpenStudy (goose.):

Well first you need to distribute that 3, so you'll get 12x-6x = 8

OpenStudy (goose.):

Combine your like terms, so 12x - 6x = 6x so 6x = 8

OpenStudy (♪chibiterasu):

You don't really need to distribute first since they are the variable. So you can subtract first.

OpenStudy (♪chibiterasu):

And you need to isolate 6, 6x isn't an answer. You divide 6 on both sides to cancel out and find the actual solution.

OpenStudy (♪chibiterasu):

\[\frac{8}{6} \approx 1.3\]

OpenStudy (♪chibiterasu):

It's a repeating decimal but if you plug it in you get 7.8.